A rock is thrown horizontally from the top of a cliff 88 m high, with a horizontal speed of 25 m/s. with what velocity does the rock hit


1
Expert's answer
2022-06-13T16:20:47-0400
"x(t)=v_{0x}t"

"y(t)=h+0\cdot t-\dfrac{gt^2}{2}"

"v_y(t)=y'(t)=-gt"

The rock hits the land: "y(t_1)=0."


"h-\dfrac{gt_1^2}{2}=0, t_1\ge0"

"t_1=\sqrt{\dfrac{2h}{g}}"

"v_y(t_1)=-g\sqrt{\dfrac{2h}{g}}=-\sqrt{2gh}"

"|v(t_1)|=\sqrt{v_{0x}^2+(-\sqrt{2gh})^2}"

"=\sqrt{v_{0x}^2+2gh}"

"=\sqrt{(25m/s)^2+2(9.81m/s^2)(88m)}"

"=48.493m\/s"

"\tan \theta=-\dfrac{\sqrt{2gh}}{v_{0x}}=-\dfrac{\sqrt{2(9.81m/s^2)(88m)}}{25m/s}"

"\approx-1.662"

"\theta\approx121\degree"

The rock hits the land with the velocity 48.493 m/s at the angle "121\degree" with respect to the ground.